1/2+1/6+1/12+1/20+1/30+....+1/n(n+1)
Answer by richard1234(7193)

Hint: 1/n(n+1) = 1/n - 1/(n+1). For example 1/6 = 1/2 - 1/3. Do that and see what you get.